Grade 11Physical Chemistry

The volume of 1.204×10^24 molecules of water at 4 cecius is

Pradeep
7 Years agoGrade 11

1 Answer

Akm
7 Years ago
Volume of 1 mole water= 1000g/l
1 mole contains avogardos no of molecules.
So volume of 1.204×10^24 molecule=( 1000÷(6.023×10^23))×1.204×10^24
=1.37×10^-45